OpenStudy (anonymous):

The vector -5.2A⃗ has a magnitude of 35m and points in the positive x direction. Find the x component of the vector A and the magnitude of the vector A⃗

OpenStudy (anonymous):

if i understood correctly your problem: \[-5.2*A_x=+35 <=> A_x=-35/5.2=-6.73 \ m \\ \left| A \right| =\sqrt{A_x^2+0+0}=6.73 \ m\]

OpenStudy (anonymous):

ohh. i see what you did. can't believe i didn't think of that. thanks!